OpenStudy (xochital):

Find the exact value of the expression 4 the square root of 13. + the square root of 6. - 7 the square root of 13..

OpenStudy (xochital):

the equation is \[4\sqrt{13} + \sqrt{6} - 7\sqrt{13}\]

OpenStudy (anonymous):

Find the reciprocal of this number.

OpenStudy (anonymous):